OpenStudy (anonymous):

Which is the set of roots of the equation? 0 = x2 + 10x - 24

OpenStudy (anonymous):

x2+12x-2x-24=0 x(x+12)-2(x+12)=0 (x-2)(x+12)=0 x-2=0 x=2 and x+12=0 x=-12

OpenStudy (anonymous):

x2+12x-2x-24=0 x(x+12)-2(x+12)=0 (x-2)(x+12)=0 (x-2)=0 or (x+12)=0 x=2 or x=-12